# Minimizing resource overhead in fusion-based quantum computation using hybrid spin-photon devices

### v1

@@ III.1.3 Transmission efficiency
In practice, reaching the upper bound on resource efficiency is challenging, as improving $\eta_{\text{mux}}$ to unity often demands a complex multiplexing network capable of flexibly redistributing photons in space and time for use by the RSG at the right place and time [51]. However, increasing the complexity of the multiplexing network reduces the transmission efficiency $\eta_{T}$ of photons in the RSG due to the greater number of optical components, such as beam splitters, phase shifters, mode crossings, and delay lines, required for photon routing. For this reason, it is useful to distinguish between the transmission efficiency that depends on the specific choice of multiplexing network, denoted here as $\eta_{x}$, and the efficiency of photons produced by the source, which we denote $\beta$. This latter efficiency impacts all types of sources and it quantifies the transmission between the source and the input of the RSG network, including the collection efficiency of the photons at the output of the source and loss due to filtering or frequency separation. This gives a total RSG transmission efficiency of
@@
Note that both $\eta_{T}$ and the RSG success probability $p_{\text{rsg}}$ impact the logical error rate of the fusion network but they do so in different ways. A non-unity $\eta_{T}$ will largely cause uncorrelated photon loss errors whereas non-unity $p_{\text{rsg}}$ results from the failure of at least one probabilistic operation inside the RSG that may introduce correlated errors. The error-correction threshold for $\eta_{T}$ depends on the resource state being produced, and is known to be $92.5\%$ [16]. However, the error-correction threshold for $p_{\text{rsg}}$ depends on how the state is constructed and would require a detailed error correction simulation of the fusion network to compute. Such a task goes beyond the scope of this comparative study and so we instead fix $p=0.999\gg 0.925$ so that any errors are dominated by independent photon loss. In addition, we verify that changing the value of $p_{\text{rsg}}$ does not significantly impact the conclusions drawn from our study.
@@ III.1.4 Per-component maximal loss
Setting aside $\beta$ for now, we analyse $\eta_{x}\geq 0.925$. This requirement is satisfied by many possible configurations of the RSG that each correspond to a different value of $\eta_{R}$. It is thus important and nontrivial to find the maximum resource efficiency such that the RSG operates in the loss-tolerant regime. For a given architecture, and considering only $D$ optical components that dominate the losses with a similar loss-per component of $\sim x$, we can write $\eta_{x}=(1-x)^{D}$. Here, we denote $D$ as the optical depth of the network, or the number of the most-lossy components seen by a photon passing through the RSG. Importantly a lower bound on $D$ can be estimated for a particular architecture, which gives a per-component maximal loss $x$ such that $\eta_{x}\geq 0.925$ using:
@@ III.2 Performance bounds
HSPS, $p_{s}$=0.005 All-photonic $6$ $36$ $0.00057\%$ $423400$ $0.22\%$
HSPS, $p_{s}$=0.05 All-photonic $6$ $36$ $0.0057\%$ $42340$ $0.22\%$
HSPS, $p_{s}$=0.25 All-photonic $6$ $36$ $0.029\%$ $8468$ $0.22\%$
DSPS All-photonic $5$ $30$ $0.11\%$ $2117$ $0.26\%$
4-GHZ source Hybrid $3$ $18$ $0.77\%$ $310$ $0.43\%$
Caterpillar source Caterpillar $2$ $12$ $8.6\%$ $28$ $0.65\%$
RUS Module RUS 00 $1$ $20\%$ $12$ $7.5\%$
Table 1: Performance bounds to generate a Shor-encoded (2,2) 6-ring resource state with the proposed architectures. For each source and generation scheme, the third column gives the number of multiplexing stages and the fourth column gives the estimated minimum optical depth associated to the multiplexing network. The maximum resource efficiency in the fifth column is given by Eq. (2), and an example of number of SPS reaching this bound is computed via Eq. 1 for a rate $r=100$MHz and displayed in the sixth column. The last column gives the maximal loss per component that each architecture can tolerate to reach a total transmission efficiency of $92.5\%$ [16] when $\beta=1$, given by Eq. 6. Since $\beta$ contributes only once to the total transmission, this maximal loss per component also estimates the required source efficiency to be $\beta\geq 1-x$. The non-RUS architectures use fusion gates boosted to a success probability of $3/4$. All sources are taken to operate at a base repetition rate of $r_{0}=1$ GHz, which is expected to be limited by recovery times for detectors and feedforward operations.
@@ III.2.1 All-photonic scheme
In the Supplementary, we formally extend this method for any number of MUX stages. This gives a recursive formula to compute the cumulative average number of photons consumed to succeed the $i$th stage:
@@
where $N_{\text{avg},0}=1$. Note that this expression gives $N_{\text{avg},3}\simeq 2\times 309$ for producing two 4-GHZ states.
The values of $p_{i}$, $c_{i}$ and $a_{i}$ for the six-stage all-photonic scheme are summarized in Table 2. Following Eq. (7) using the values in the table, we arrive at the total of $N_{\text{avg}}=21170$ single photons for one resource state. Finally, from Eq. (1), we have the upper bound
@@
For a heralded SPS with $p_{s}=0.5\%$ success probability, the resource efficiency is then bounded by $\eta_{R}\leq 0.00057\%$, while for a deterministic source it is $\eta_{R}\leq 0.11\%$. Assuming that the internal clock cycle (the pump repetition rate) is $r_{0}=1$ GHz and we wish to produce resource states at the rate $r=100$ MHz, these resource efficiencies correspond to a minimum number of sources of $423400$ and $2117$, respectively.
@@
Probability, $p_{i}$ $1$ $2820\sqrt{2}-$ 28 $\dfrac{3+2\sqrt{2}}{64}$ $\dfrac{9}{16}$ $\dfrac{9}{16}$ $\dfrac{9}{16}$
@@
Cumul. $N_{\text{avg},i}$ $4$ $28$ $618$ $3339$ $11900$ $21170$
Cumulative $D_{i}$ $9$ $15$ $21$ $27$ $33$ $36$
Table 2: A summary of the relevant quantities describing the multiplexing stages of the proposed all-photonic scheme to building a Shor encoded (2,2) 6-ring resource state. Here, $p_{i}$ is the success probability of the $i$th stage operation in the absence of loss, $c_{i}$ is the number of copies needed of the state produced in that stage to attempt the following stage, $a_{i}$ is the number of auxiliary single photons needed for boosting fusion gates, $d_{i}$ is the number of detected photons, and  $N_{\text{avg, $i$}}$ is the cumulative average number of single photons needed to succeed the stage. $D_{i}$ is the cumulative optical depth needed to construct the resource state and to route the photons through the network.
To estimate the maximal loss per component $x$ required to remain in the fault-tolerant regime, we first make an optimistic approximation that the optical depth of a photonic MUX does not depend on the number of photonic states being routed between each stage. That is, we assume that the optical depth of the RSG depends only on the type of MUX networks used and the number of MUX stages. For the type of spatial MUX network, we consider a Spanke network that uses two layers of generalized Mach-Zehnder interferometers (GMZIs) with a mode routing network in between [51]. Each GMZI is composed of many beam splitters, but each photon only sees one phase shifter using the most optimistic design [51]. Since a phase shifter is expected to be the most lossy component for a GMZI, we estimate the minimum optical depth to be $D_{\text{GMZI}}=1$. The mode routing network requires photons to be redistributed in space, which requires long propagation distances and for many lossy mode crossings. For this reason, the practical solution is to couple photons into low-loss fibers to redistribute them arbitrarily with minimal propagation loss. However, each chip-to-fiber interface causes non-negligible loss, which we identify as the most lossy component. Since it is necessary to couple out of the first GMZI and into the second GMZI, we estimate the minimum constant optical depth of the mode routing network to be $22D_{\text{coupl.}}=$ 2.
@@ III.2.2 Caterpillar scheme
Detected, $d_{i}$ 0 11 12
Cumulative $N_{\text{avg},i}$ $17$ $112$ $280$
@@
Table 3: A summary of the relevant quantities describing the multiplexing stages of the proposed caterpillar scheme to building a Shor encoded (2,2) 6-ring resource state. Here, $p_{i}$ is the success probability of the $i$th stage operation in the absence of loss, $c_{i}$ is the number of copies needed of the state produced in that stage to attempt the following stage, $a_{i}$ is the number of auxiliary single photons needed for boosting fusion gates, $d_{i}$ is the number of detected photons, and $N_{\text{avg,$i$}}$ is the cumulative average number of single photons needed to succeed the stage. $D_{i}$ is the cumulative optical depth needed to construct the resource state and to route the photons through the network.
For caterpillar scheme, we can compute the average number of photons $N_{\text{avg}}$ required to construct a single resource state using the same formula as for the all-photonic scheme. The probabilities, number of copies, and auxiliary photons used for fusion gates are all summarized in Table 3. Applying Eq. 7, we find that the average number of photons consumed to produce one resource state on average is $N_{\text{avg}}=280$. This corresponds to an upper bound on the resource efficiency of $\tilde{\eta}_{R}=8.6\%$ for the $N=24$-photon state.
@@ III.2.3 RUS scheme
Each gate attempt succeeds with probability $1/2$ [18, 8, 9], therefore each round of RUS gates requires 2 attempts on average. Since the maximum degree of the LC-graph is 4 (recall Figure 3c), 4 rounds of RUS gates are necessary to produce the spin graph. Thus, all 12 sources must wait for an average of 8 RUS attempts before the LC-graph is complete. Given that local gates do not require feedforward, we assume they are completed much faster than the RUS gates so that they have a negligible impact on $r$. Finally, we account for two emission cycles that we assume operate at the same rate $r_{0}$ for simplicity. In total, the resource state can be generated using on average $\tau_{\text{avg}}=10$ internal clock cycles. This gives the RUS module approach a maximum resource efficiency of $\tilde{\eta}_{R}=24/(12\times 10)=20\%$.
As discussed in Section II.2.3, the RUS scheme was designed to minimize the optical depth to a single $1\times 2$ switch. In addition, although some platforms may use chip-to-fiber coupling to route photons through the switch, this is not strictly necessary because photons are stored in the spin memory until the switch is reconfigured. A single switch is also significantly more shallow than a GMZI. For these reasons, we estimate a minimum optical depth of $D=1$ for the RUS scheme. This implies that the per-component maximal loss is $x=7.5\%$, which is significantly higher than all other architectures thanks to the lossless memory of the RUS module. Note that, since $7.5\%$ is also substantially higher than the loss from a $1\times 2$ switch, the loss will be dominated by the source itself, meaning that a source with $\beta\geq 92.5\%$ reaches the loss-tolerant regime for this architecture.
@@ III.3 Implementations
Figure 4: Resource efficiency and maximal loss that each architecture can tolerate. a. The regimes where fault-tolerant quantum computing is possible using the 24-photon Shor-encoded (2, 2) 6-ring resource state are indicated by shaded areas. The square points represent the hypothetical optimal implementations presented in Table 1. The curves illustrate the simulated trade-off between resource efficiency and maximal loss per component when using resource-sharing MUX strategies. For the HSPS, we assume a photon generation success probability of $p_{s}=5\%$, representing realistic improvements over current technology. For the RUS scheme, the solid line represents a single 12-source RSG while the dashed line represents the maximum resource efficiency achieved when using spatial multiplexing to reach $r=r_{0}$. b. The number of sources needed to reach an RSG clock rate of $r=1$ GHZ assuming an internal clock rate of $r_{0}=1$ GHZ and a maximal loss per component of $x=0.1\%$ for the best (RUS) and worst (HSPS) architectures shown in panel a. In both plots, we define the hypothetical ideal RSG architecture represented by the star point as an RSG that is able to produce resource states with perfect resource efficiency ($\eta_{R}=1$) and without photons passing through any optical components after being collected from source ($\eta_{x}=1$) so that the loss is purely determined by the source collection efficiency $\beta$.
@@
To illustrate these trade-offs and provide a more practical comparison of the schemes, we simulate the resource efficiency while incorporating resource-sharing strategies.
To simulate photonic multiplexing used by the all-photonic and caterpillar schemes, we assume that the RSG is composed of a series of independent RSG modules that each use maximal resource sharing at all but the last stage. These independent RSGs are then multiplexed at the final stage to achieve the desired success probability of $p_{s}=99.9\%$. For simplicity, we assume this final multiplexing stage can be accomplished in constant depth $3$ as before. The number of sources within each independent module and the resulting success probability is determined by the remaining available optical depth $D$ so that Eq. (6) is satisfied. For each MUX inside each module, we assume an optimistic sorting network of depth $\lceil\log_{2}(n)\rceil$ where $n$ is the number of states routed at the given stage. More details about the different figures and the optimization procedure are given in Supplementary. Note that less optimal sorting might be achieved with a smaller depth, but this would not allow for maximal resource sharing.
For the RUS scheme, we performed a Monte Carlo simulation of the construction procedure that is illustrated in Figure 3c (see Supplementary for details). For this scheme, we also consider two types of resource sharing. The first type is a temporal resource sharing technique that naturally occurs due to the spin memory. If a graph was not complete in time for the RSG cycle, the partially completed graph will be carried over into the next cycle, greatly increasing the probability of success. The second type is a spatial resource sharing technique that is possible due to having identical subgraphs in each stage of construction. In this case, the next stage of graph construction can commence as soon as any two subgraphs are complete, provided that the RUS gate connectivity can be reconfigured. To allow for this additional connectivity, and knowing that no delay lines are required to reconfigure the network due to the quantum memory, we assume a constant optical depth 3 on the RUS gate. Note that this does not increase the optical depth impacting the final photonic resource state, which remains $D=1$ due to the single switch.
@@
In Figure 4a, we show how four architectures compare in this practical setting using resource sharing techniques. Notably, in this practical setting, all four architectures fail to simultaneously reach the bounds on resource efficiency and maximal loss per component presented in Table 1. However, the discrepancy is much greater for the all-photonic architectures compared to the caterpillar and RUS schemes. There is also a substantial drop in resource efficiency for the all-photonic architectures beyond a certain value of loss, which is caused by a rapid reduction in resource sharing in order to reduce the optical depth of each MUX stage to remain below the loss threshold of the resource state. On the other hand, the caterpillar scheme does not suffer a large trade off because only one MUX stage uses resource sharing and it is already very shallow due to the relatively high success probability of performing two boosted fusion gates.
It is notable that the RUS scheme also suffers from a trade-off between the resource efficiency and the maximal loss per component, with a marked decrease in resource efficiency as the maximal loss per component approaches the upper bound. This is due to a reduced efficiency of the RUS gate, causing the module to restart some operations whenever a photon is lost. This increases the amount of internal clock cycles needed to ensure $p_{\text{rsg}}$ reaches 99.9%. However, when using spatial multiplexing, the resource efficiency can approach very close to the 20% upper bound for the scheme even in the presence of significant loss. This latter result is a consequence of the fact that, for this scheme uniquely, resource sharing and multiplexing primarily impact photons used to entangle sources, rather than the final resource state itself.
Our results also show that, for a maximal loss per component of $x=99.9\%$, a multiplexed RUS module requires 6 orders of magnitude fewer sources to achieve the same rate of generation as an HSPS with $p_{s}=5\%$ (see Figure 4b). Improving $p_{s}$ to at most $25\%$ can reduce the resource overhead of the all-photonic architecture by factor of around 5, but this surprisingly provides very little improvement to the loss tolerance. Moreover, due to the shallow optical depth, the RUS scheme can achieve this resource advantage using sources with an order of magnitude higher collection loss.

### v2

@@ III.1.3 Transmission efficiency
In practice, reaching the upper bound on resource efficiency is challenging, as improving $\eta_{\text{mux}}$ to unity often demands a complex multiplexing network capable of flexibly redistributing photons in space and time for use by the RSG at the right place and time [57]. However, increasing the complexity of the multiplexing network reduces the transmission efficiency $\eta_{T}$ of photons in the RSG due to the greater number of optical components, such as beam splitters, phase shifters, mode crossings, and delay lines, required for photon routing. For this reason, it is useful to distinguish between the transmission efficiency that depends on the specific choice of multiplexing network, denoted here as $\eta_{x}$, and the transmission efficiency of photons produced by the source, which we denote $\beta$. Unlike $p_{s}$, which is non-unity only for HSPSs, $\beta$ impacts all types of sources and it quantifies the transmission between the source and the input of the RSG network, including the collection efficiency of the photons at the output of the source and loss due to filtering or frequency separation. This gives a total RSG transmission efficiency of
@@
Note that both $\eta_{T}$ and the RSG success probability $p_{\mathrm{rsg}}$ impact the logical error rate of the fusion network but they do so in different ways. A non-unity $\eta_{T}$ will largely cause uncorrelated photon loss errors whereas non-unity $p_{\mathrm{rsg}}$ results from the failure of at least one probabilistic operation inside the RSG that may introduce correlated errors. The error-correction threshold for $\eta_{T}$ depends on the resource state being produced, and is known to be $92.5\%$ [16]. However, the error-correction threshold for $p_{\mathrm{rsg}}$ depends on how the state is constructed and would require a detailed error correction simulation of the fusion network to compute. Such a task goes beyond the scope of this comparative study and so we instead ensure $p_{\mathrm{rsg}}\geq 0.999\gg 0.925$ so that we explore a regime where errors are dominated by independent photon loss. In principle, FBQC with an appropriate decoding strategy could tolerate values of $p_{\mathrm{rsg}}$ lower than $0.999$, which could help increase $\eta_{R}$. However, as discussed in Appendix I, the gain in $\eta_{R}$ is bounded: relaxing $p_{\mathrm{rsg}}$ from 0.999 to 0.9 increases $\eta_{R}$ by at most a factor of 3, and typically much less if $\eta_{R}$ is already close to $\tilde{\eta}_{R}$.
@@ III.1.4 Per-component maximal loss
Setting aside $\beta$ for now, we analyze $\eta_{x}\geq 0.925$. This requirement is satisfied by many possible configurations of the RSG that each correspond to a different value of $\eta_{R}$. It is thus important and nontrivial to find the maximum resource efficiency such that the RSG operates in the loss-tolerant regime. For a given architecture, and considering only $D$ optical components that dominate the losses with a similar loss-per component of $\sim x$, we can write $\eta_{x}=(1-x)^{D}$. Here, we denote $D$ as the optical depth of the network, or the number of the most-lossy components seen by a photon passing through the entire network to be released by the RSG as part of the final resource state. Importantly, we estimate in Section III.2 below a lower bound on $D$ for each architecture, which will imply a maximal per-component loss $x$ such that $\eta_{x}\geq 0.925$ using:
@@ III.2 Performance bounds
HSPS, $p_{s}$=0.005 All-photonic 66 $36$ $0.00057\%$ $423400$ $0.22\%$
HSPS, $p_{s}$=0.05 All-photonic 66 $36$ $0.0057\%$ $42340$ $0.22\%$
HSPS, $p_{s}$=0.25 All-photonic 66 $36$ $0.029\%$ $8468$ $0.22\%$
DSPS All-photonic 55 $30$ $0.11\%$ $2117$ $0.26\%$
4-GHZ source Hybrid 33 $18$ $0.77\%$ $310$ $0.43\%$
Caterpillar source Caterpillar 22 $12$ $9.9\%$ $25$ $0.65\%$
RUS Module RUS 0 11 $20\%$ $12$ $7.5\%$
Table 1: Performance bounds to generate a Shor-encoded (2,2) 6-ring resource state with the proposed architectures. For each source and generation scheme, the third column gives the number of multiplexing stages and the fourth column gives the estimated minimum optical depth associated to the multiplexing network. The maximum resource efficiency in the fifth column corresponds to Eq. (2), except for the RUS module, for which the bound is given by Eq. (4), and an example of number of SPS reaching this bound is computed via Eq. (1) for a rate r=100r=100MHz and displayed in the sixth column. The last column gives the maximal loss per component that each architecture can tolerate to reach a total transmission efficiency of $92.5\%$ [16] when $\beta=1$, given by Eq. (6). Since $\beta$ contributes only once to the total transmission, this maximal loss per component also estimates the required source efficiency to be $\beta\geq 1-x$. The non-RUS architectures use fusion gates boosted to a success probability of $3/4$. All sources are taken to operate at a base repetition rate of $r_{0}=1$ GHz, which is expected to be limited by recovery times for detectors and feedforward operations.
@@ III.2.1 All-photonic scheme
In Appendix D, we extend this method to any number of MUX stages. This gives a recursive formula to compute the cumulative average number of photons consumed to succeed the $i$th stage:
@@
where $N_{\text{avg},0}=1$. Similar recursive formulas have been derived in previous work to estimate resource overhead for photonic state construction [55].
The values of $p_{i}$, $c_{i}$ and $a_{i}$ for the six-stage all-photonic scheme are summarized in Table 2. Note that standard fusion gates succeed with a probability of $50\%$. Here, to improve the probability of successfully generating the state, we assume that each gate is boosted to reach a success probability of $75\%$ using four additional single photons (not shown in Fig. 3a) [46]. Indeed, in the low-loss regime considered in our study, the cost of producing additional photons is lower than the cost of multiplexing more fusion gates. However, boosting beyond $75\%$ requires additional GHZ states, which are themselves costly to produce from single photons and do not justify the mild increase to success probability compared to increasing the amount of multiplexing.
Following Eq. (7) using the values in the table, we arrive at the total of $N_{\text{avg}}=21170$ single photons for one resource state. Finally, from Eq. (1), we have the upper bound
@@
For a heralded SPS with $p_{s}=0.5\%$ success probability, the resource efficiency is then bounded by $\eta_{R}\leq 0.00057\%$, while for a deterministic source it is $\eta_{R}\leq 0.11\%$. Assuming that the internal clock cycle (the pump repetition rate) is $r_{0}=1$ GHz and we wish to produce resource states at the rate $r=100$ MHz, these resource efficiencies correspond to a minimum number of sources of $N_{0}=423400$ and $N_{0}=2117$, for the heralded and deterministic SPSs respectively.
@@
Probability, $p_{i}$ 11 $20\sqrt{2}-28$ $\dfrac{3+2\sqrt{2}}{64}$ $\dfrac{9}{16}$ $\dfrac{9}{16}$ $\dfrac{9}{16}$
@@
Cumul. $N_{\text{avg},i}$ 44 $28$ $618$ $3339$ $11900$ $21170$
Cumulative $D_{i}$ 99 $15$ 2121 $27$ $33$ $36$
Table 2: A summary of the relevant quantities describing the multiplexing stages of the proposed all-photonic scheme to building a Shor encoded (2,2) 6-ring resource state. Here, $p_{i}$ is the success probability of the $i$th stage operation in the absence of loss, $c_{i}$ is the number of copies needed of the state produced in that stage to attempt the following stage, $a_{i}$ is the number of auxiliary single photons needed to boost fusion gates using Ewert and van Loock’s scheme [61], $d_{i}$ is the number of detected photons, and $N_{\text{avg, $i$}}$ is the cumulative average number of single photons needed to succeed the stage. $D_{i}$ is the cumulative optical depth needed to construct the resource state and to route the photons through the network.
To estimate the maximal loss per component $x$ required to remain in the fault-tolerant regime, we first make an optimistic approximation that the optical depth of a photonic MUX does not depend on the number of photonic states being routed between each stage. That is, we assume that the optical depth of the RSG depends only on the type of MUX networks used and the number of MUX stages. For the type of spatial MUX network, we consider a Spanke network that uses two layers of generalized Mach-Zehnder interferometers (GMZIs) with a mode routing network in between [57]. Each GMZI is composed of many beam splitters, but each photon only sees one phase shifter using the most optimistic design [57]. Since a phase shifter is expected to be the most lossy component for a GMZI, we estimate the minimum optical depth to be $D_{\text{GMZI}}=1$. The mode routing network requires photons to be redistributed in space, which requires long propagation distances and for many lossy mode crossings. For this reason, the practical solution is to couple photons into low-loss fibers to redistribute them arbitrarily with minimal propagation loss. However, each chip-to-fiber interface causes non-negligible loss, which we identify as the most lossy component. Since it is necessary to couple out of the first GMZI and into the second GMZI, we estimate the minimum constant optical depth of the mode routing network to be $2D_{\text{coupl.}}=2$.
@@ III.2.2 Caterpillar scheme
Detected, $d_{i}$ 0 8 12
Cumulative $N_{\text{avg},i}$ 14 96 242
@@
Table 3: A summary of the relevant quantities describing the multiplexing stages of the proposed caterpillar scheme to building a Shor encoded (2,2) 6-ring resource state. Here, $p_{i}$ is the success probability of the $i$th stage operation in the absence of loss, $c_{i}$ is the number of copies needed of the state produced in that stage to attempt the following stage, $a_{i}$ is the number of auxiliary photons needed to boost fusion gates using Grice’s scheme [40], $d_{i}$ is the number of detected photons, and $N_{\text{avg,$i$}}$ is the cumulative average number of single photons needed to succeed the stage. $D_{i}$ is the cumulative optical depth needed to construct the resource state and to route the photons through the network.
For the caterpillar scheme, we can compute the average number of photons $N_{\text{avg}}$ required to construct a single resource state using the same formula as for the all-photonic scheme. The probabilities, number of copies, and auxiliary photons used for fusion gates are all summarized in Table 3. Here, we assume that each fusion is again boosted to a 75% success probability, but this time using an auxiliary Bell pair [40]. Note that a caterpillar source can also be used to efficiently boost fusion gates beyond 75%, however we find that the relative improvement compared to once-boosted gates is minor (see Appendix E) and may not justify the additional losses due to the increased complexity. Hence, we focus on the once-boosted fusion gate in our analysis.
Applying Eq. 7, we find that the average number of photons consumed to produce one resource state on average is $N_{\text{avg}}=242$. This corresponds to an upper bound on the resource efficiency of $\tilde{\eta}_{R}=9.9\%$ for the $N=24$-photon state.
@@ III.2.3 RUS scheme
Each gate attempt succeeds with probability $1/2$ [18, 8, 9], therefore each round of RUS gates requires 2 attempts on average. Since the maximum degree of the LC-graph is 4 (recall Figure 3c), 4 rounds of RUS gates are necessary to produce the spin graph. Thus, all 12 sources must wait for an average of 8 RUS attempts before the spin graph is complete. Finally, we account for two emission cycles that we assume operate at the same rate $r_{0}$ for simplicity. In total, the resource state can be generated using on average $\tau_{\text{avg}}=10$ internal clock cycles. Using Eq. (4), this gives the RUS module approach a maximum resource efficiency of $\tilde{\eta}_{R}=24/(12\times 10)=20\%$. Note that this bound neglects the time needed to implement local operations such as single-qubit gates and measurements. In practice, these operations can be implemented much faster than the feedforward-limited RUS gates, and so we assume they have a negligible impact on $\tau_{\text{avg}}$ (see Appendix H for more discussion on operation timescales).
As discussed in Section II.2.3, the RUS scheme was designed to minimize the optical depth to a single $1\times 2$ switch. In addition, although some platforms may use chip-to-fiber coupling to route photons through the switch, this is not strictly necessary because photons are stored in the spin memory until the switch is reconfigured. A single switch is also significantly more shallow than a multiplexing stage needed to route probabilistic inputs. For these reasons, we estimate a minimum optical depth of $D=1$ for the RUS scheme. This implies that the per-component maximal loss is $x=7.5\%$, which is significantly higher than all other architectures thanks to the lossless memory of the RUS module. Note that, since $7.5\%$ is also substantially higher than the anticipated loss from a single active switch, the loss will be dominated by the source itself, meaning that a source with $\beta\geq 92.5\%$ reaches the loss-tolerant regime for this architecture.
@@ III.3 Implementations
Figure 4: Resource efficiency and maximal loss that each architecture can tolerate. a. The regimes where fault-tolerant quantum computing is possible using the 24-photon Shor-encoded (2, 2) 6-ring resource state are indicated by shaded areas. The square points represent the hypothetical optimal implementations presented in Table 1. The curves illustrate the simulated trade-off between resource efficiency and maximal loss per component when using MUX strategies to reach RSG success probability $p_{\mathrm{rsg}}=99.9\%$. For the HSPS, we assume a photon generation success probability of $p_{s}=5\%$, representing realistic improvements over current technology. For the RUS scheme, the solid line represents a single 12-source RSG while the dashed line represents the maximum resource efficiency achieved when using spatial multiplexing to reach $r=r_{0}$ at the cost of increased depth $D=2$. Both RUS curves consider only temporal resource sharing. b. The number of sources needed to reach $p_{\mathrm{rsg}}=99.9\%$ at $r=r_{0}=1$ GHz for the best (RUS) and worst (HSPS) architectures shown in panel a using resource sharing, spatial multiplexing, and a maximal loss per component of $x=0.1\%$. In both plots, we define the hypothetical ideal RSG architecture represented by the star point as an RSG that is able to produce resource states with perfect resource efficiency ($\eta_{R}=1$) and without photons passing through any optical components after being collected from source ($\eta_{x}=1$) so that the loss is purely determined by the collection efficiency $\beta$.
@@
To illustrate these trade-offs and provide a more practical comparison of the schemes, we simulate the resource efficiency while incorporating realistic multiplexing and resource-sharing strategies.
To simulate photonic multiplexing used by the all-photonic and caterpillar schemes, we assume that the RSG is composed of a series of independent RSG modules that each use maximal resource sharing at all but the last stage. These independent RSGs are then multiplexed at the final stage to achieve the desired success probability of $p_{\mathrm{rsg}}=99.9\%$. For simplicity, we assume this final multiplexing stage can be accomplished in constant depth $3$ as before. The number of sources within each independent module and the resulting success probability is determined by the remaining available optical depth $D$ so that Eq. (6) is satisfied. For each MUX inside each module, we assume an optimistic sorting network of depth $\lceil\log_{2}(n)\rceil$ where $n$ is the number of states routed at the given stage. More details about the different figures and the optimization procedure are given in Appendix G. Note that less optimal sorting might be achieved with a smaller depth, but this would not allow for maximal resource sharing.
For the RUS scheme, we performed a Monte Carlo simulation of the construction procedure that is illustrated in Figure 3c (see Appendix H for details). For this scheme, we also consider two types of resource sharing. The first type is a temporal resource sharing technique that naturally occurs due to the spin memory. If a graph was not complete in time for the RSG cycle, the partially completed graph will be carried over into the next cycle, greatly increasing the probability of success. The second type is a spatial resource sharing technique that is possible due to having identical subgraphs in each stage of construction. In this case, the next stage of graph construction can commence as soon as any two subgraphs are complete, provided that the RUS gate connectivity can be reconfigured. To allow for this additional connectivity, and knowing that no delay lines are required to reconfigure the network due to the quantum memory, we assume a constant optical depth 3 on the RUS gate when using spatial resource sharing. Note that this does not increase the optical depth impacting the final photonic resource state, which remains $D=1$ due to the single switch.
@@
In Figure 4a, we show how four architectures compare in this practical setting using multiplexing and resource sharing techniques. Notably, in this practical setting, all four architectures fail to simultaneously reach the bounds on resource efficiency and maximal loss per component presented in Table 1. However, the discrepancy is much greater for the all-photonic architectures compared to the caterpillar and RUS schemes. There is also a substantial drop in resource efficiency for the all-photonic architectures beyond a certain value of loss, which is caused by a rapid reduction in resource sharing in order to reduce the optical depth of each MUX stage to remain below the loss threshold of the resource state. On the other hand, the caterpillar scheme does not suffer a large trade-off because only one MUX stage uses resource sharing and it is already very shallow due to the relatively high success probability of performing two boosted fusion gates.
It is notable that the RUS scheme also suffers from a trade-off between the resource efficiency and the maximal loss per component, with a marked decrease in resource efficiency as the maximal loss per component approaches the upper bound. This is due to a reduced efficiency of the RUS gate, causing the module to restart some operations whenever a photon is lost (see Appendix H for more information on erasure errors). This increases the amount of internal clock cycles needed to ensure $p_{\mathrm{rsg}}$ reaches 99.9%. However, when using spatial multiplexing, the resource efficiency can approach very close to the 20% upper bound for the scheme even in the presence of significant loss. This latter result is a consequence of the fact that, for this scheme uniquely, resource sharing and multiplexing primarily impact photons used to entangle sources, rather than the final resource state itself.
Our results also show that, for a maximal loss per component of $x=0.1\%$ and when considering spatial resource sharing, a multiplexed RUS module requires nearly six orders of magnitude fewer sources to achieve the same rate of generation as an HSPS with $p_{s}=5\%$ (see Figure 4b). Improving $p_{s}$ to at most $25\%$ can reduce the resource overhead of the all-photonic architecture by factor of around five, but this surprisingly provides very little improvement to the loss tolerance. Moreover, due to the shallow optical depth, the RUS scheme can achieve this resource advantage using sources with an order of magnitude higher collection loss.